# Flexible game-playing AI with AlphaViT: adapting to multiple games and board sizes

### v1

@@ Abstract
This paper presents novel game AI agents based on the AlphaZero framework, enhanced with Vision Transformers (ViT): AlphaViT, AlphaViD, and AlphaVDA. These agents are designed to play various board games of different sizes using a single model, overcoming AlphaZero’s limitation of being restricted to a fixed board size. AlphaViT uses only a transformer encoder, while AlphaViD and AlphaVDA contain both an encoder and a decoder. AlphaViD’s decoder receives input from the encoder output, while AlphaVDA uses a learnable matrix as decoder input. Using the AlphaZero framework, the three proposed methods demonstrate their versatility in different game environments, including Connect4, Gomoku, and Othello. Experimental results show that these agents, whether trained on a single game or on multiple games simultaneously, consistently outperform traditional algorithms such as Minimax and Monte Carlo tree search using a single DNN with shared weights, while approaching the performance of AlphaZero. In particular, AlphaViT and AlphaViD show strong performance across games, with AlphaViD benefiting from an additional decoder layer that enhances its ability to adapt to different action spaces and board sizes. These results may suggest the potential of transformer-based architectures to develop more flexible and robust game AI agents capable of excelling in multiple games and dynamic environments.
@@ 1 Introduction
In recent years, artificial intelligence (AI) has made remarkable progress, demonstrating its potential in a wide range of applications. One application area where AI has shown significant prowess is in mastering board games, surpassing the skills of top human players in many games. Historical achievements include AI outperforming humans in games such as checkers, chess (Campbell et al., 2002), and Othello (Buro, 1997). A significant milestone was reached in 2016 when AlphaGo (Silver et al., 2016), an AI specialized in the game of Go, defeated one of the world’s top players. The subsequent introduction of AlphaZero (Silver et al., 2018), capable of mastering multiple board games such as Chess, Shogi, and Go, has further solidified the superhuman capabilities of AI in this domain.
@@
However, current game-playing AI agents have a significant drawback: they are designed to play only one specific game and cannot play other games. Even if the rules remain the same, they cannot handle variations in board size. In contrast, humans can easily switch between different board sizes. For instance, Go beginners often start by practicing on smaller boards (e.g., $9\times 9$) before moving on to larger boards (e.g., $19\times 19$). However, AI agents such as AlphaZero, which are designed for a specific game and fixed board size, cannot adapt to these changes without significant reprogramming.
@@
For AlphaZero specifically, the core of this limitation lies in the architecture of AlphaZero’s deep neural network (DNN), which requires a fixed input size. AlphaZero uses a DNN consisting of residual blocks and multilayer perceptrons (MLPs). These components are designed for a fixed input size. The output size of the residual blocks varies as the input size changes, creating an inconsistency with the expected output size for MLPs. As a result, AlphaZero cannot function properly when faced with even small variations in board size. To address this issue, we propose to replace the residual blocks in the AlphaZero framework with Vision Transformers (ViT) (Dosovitskiy et al., 2021).
ViT is an image classification DNN based on the transformer architecture. ViT divides an image into several patches and infers a class of images from these patches. A key advantage of ViT is its ability to process images independently of input size, allowing for flexible adaptation to different board sizes within the AlphaZero framework.
@@
In this study, we present game-playing agents that use a single DNN to handle multiple games and variable board sizes. These agents, named AlphaViT, AlphaViD (AlphaViT with a transformer decoder layer), and AlphaVDA (AlphaViD with learnable action tokens), are based on the AlphaZero framework. The agents predict the value and policy of game states using a DNN, while decisions are made using Monte Carlo Tree Search (MCTS). Our computational experiments show that these models can simultaneously play games such as Connect4, Gomoku, and Othello using a single DNN with shared weights. Moreover, the proposed agents outperform traditional algorithms such as Minimax and MCTS in various games, while approaching the performance of AlphaZero, whether trained on a single game or on multiple games simultaneously.
@@ 2 Related work
Game-playing AI agents have achieved superhuman-level performance in traditional board games such as Checkers (Schaeffer et al., 1993), Othello (Buro, 1997, 2003), and Chess (Campbell, 1999; Hsu, 1999; Campbell et al., 2002). In 2016, AlphaGo (Silver et al., 2016), a Go-playing AI, has defeated the world’s top Go player, becoming the first superhuman-level Go-playing AI. AlphaGo relies on supervised learning from a large database of expert human moves and self-play data. Subsequently, AlphaGo Zero (Silver et al., 2017) has defeated AlphaGo without preparing a large training dataset. In 2018, Silver et al. (2018) have proposed AlphaZero, which has no restrictions on playable games. AlphaZero has outperformed other superhuman-level AIs at Go, Shogi, and Chess. Interestingly, AlphaZero’s versatility extends beyond traditional two-player perfect information games. Research has explored its potential in more complex scenarios. For example, Hsueh et al. (2018) have shown AlphaZero’s potential in nondeterministic games. Other extensions include handling continuous action spaces (Moerland et al., 2018) and support for multiplayer games (Petosa and Balch, 2019). However, AlphaZero cannot play various games simultaneously or handle games with the same rules but different board sizes using a single DNN with shared weights.
This limitation is due to the use of the DNN in AlphaZero for policy and value estimation. AlphaZero’s DNN consists of residual blocks and MLPs. While the residual blocks excel at extracting features from input images, the MLPs are constrained by a fixed input size. As a result, the DNN is not sufficiently flexible to accommodate variations in board size. To address this limitation, the integration of Vision Transformer (ViT) (Dosovitskiy et al., 2021) into AlphaZero may provide a solution.
@@
The transformer architecture, initially designed for natural language processing (Vaswani et al., 2017), has shown remarkable effectiveness in various domains. The transformer architecture has also been successfully applied to image-processing tasks. Transformer-based models have achieved exceptional performance in various image-related tasks, including image classification (Dosovitskiy et al., 2021), semantic segmentation (Xie et al., 2024), video classification (Li et al., 2022), and video captioning (Zhao et al., 2022). ViT, introduced by Dosovitskiy et al. (Dosovitskiy et al., 2021), is a remarkable example of a transformer-based model for image processing. ViT has achieved state-of-the-art performance in image classification at the time of its introduction.
A key feature of ViT is its independence from the size of the input image (Dosovitskiy et al., 2021). Unlike convolutional neural networks (CNNs), which require fixed-size inputs, ViT can handle images of various sizes by dividing them into fixed-size patches and treating each patch as a token in the transformer architecture. This flexibility allows ViT to be highly adaptable and efficient in handling different image sizes.
While AlphaZero can only play the game it was trained on, humans can play multiple games with a single brain, such as Chess, Othello, and Connect4. In addition, humans can adapt to different board sizes if they know the rules of a game. However, even if only the board size of a game changes, AI such as AlphaZero cannot play the game. The goal of this study is to overcome this limitation by developing AI agents based on AlphaZero, which can be generalized across different board sizes and various games.
@@ 3 AlphaViT, AlphaViD, and AlphaVDA
AlphaZero’s game-playing capability is constrained to games with identical board sizes and rules as those used during its training. This limitation is due to AlphaZero’s ResNet-based deep neural network (DNN), which consists of residual blocks followed by multilayer perceptrons (MLPs) for value and policy computation. The MLPs receive input from the final residual block, assuming a fixed input size. Consequently, AlphaZero’s DNN cannot adapt to variations in board size. Therefore, AlphaZero’s performance is limited to games with fixed board size, restricting its performance to games with predetermined dimensions.
@@
To overcome this limitation, AlphaViT, AlphaViD, and AlphaVDA have been developed as game-playing AIs based on AlphaZero but using Vision Transformer (ViT) architecture. These game-playing AI agents use a combination of a DNN and MCTS (Fig. 1). The DNN receives board states and outputs value estimates and move probabilities (policies). The MCTS then searches the game tree using these outputs. The MCTS searches a game tree using the estimated value and move probability. By incorporating ViT instead of residual blocks, AlphaViT, AlphaViD, and AlphaVDA can overcome the limitation of AlphaZero and can play games with different board sizes and rules.
An overview of AlphaViT’s DNN architecture is shown in Fig. 2. The DNN of AlphaViT is based on ViT, which has no image-size limitation and can classify an image even if the input image size differs from the training image size. This flexibility allows AlphaViT and AlphaViD to play games with different board sizes using the same network.
@@
In AlphaViT, the boards are input to ViT. Initially, these inputs are transformed into patch embeddings using a convolutional layer. Using the convolution layer allows for easy adjustment of patch division parameters such as patch size, stride, and padding. The final outputs of the encoder layer are then used to compute value and policy estimates.
@@
AlphaViD and AlphaVDA incorporate both transformer encoder and decoder layers for value and move probability calculations. In AlphaViD, the decoder layer receives input derived from the encoder layer’s output. Conversely, AlphaVDA utilizes learnable embeddings as input for its decoder layer.
@@
Importantly, AlphaViT, AlphaViD, and AlphaVDA can play any game that AlphaZero can, as they employ the same fundamental game-playing algorithm. Their enhanced flexibility in handling various board sizes and game rules represents a significant advancement in AI game-playing capabilities.
@@
Figure 1: AlphaViT and AlphaViD use the same procedure as AlphaZero. They receive board states and a current player, and decide a move using MCTS. MCTS searches a game tree using the value and policy estimated by the DNN.
@@
Figure 2: Overview of the AlphaViT model architecture. A board state is divided into fixed-size patches and linearly embedded by a convolutional layer. Position embeddings are added to these patch embeddings. The resulting sequence of patch embeddings is appended with the value embedding and the game embedding. The sequence is fed to a transformer encoder. Finally, the value and the move probability are estimated using a multilayer perceptron (MLP) that takes the outputs from the last layer of the transformer encoder as input.
@@ 3.1 Architectures of DNNs
AlphaViT’s deep neural network (DNN) predicts the value $v(s)$ and the move probability distribution $\bm{p}$ with components $p(a|s)$ for each action $a$, given a game state $s$. In a board game, $s$ represents the state of a board, and $a$ denotes a move.
@@
The input to the DNN is an $H\times W\times(2T+1)$ image stack ${\rm\bm{x}}\in R^{H\times W\times(2T+1)}$, consisting consisting of $12T+$ 1 binary feature planes of size $H\times W$. Here, $H$ and $W$ are the dimensions of the board, and $T$ is the number of history planes. The first $T$ feature planes represent the occupancies of the first player’s discs, where a feature value of 1 indicates that a disc occupies the corresponding cell and 0 otherwise. The following $T$ feature planes represent the occupancies of the second player’s discs. The final feature level represents the disc color of the current player, where 1 and -1 represent the first and second players, respectively.
@@
The convolutional layer processes the image stack ${\rm\bm{x}}$ to generate the patch embeddings $x_{p}$. This layer divides the image stack into $P\times P$ image patches with stride $s$ and padding $p$ and reshapes them into flattened 2D patches ${\rm\bm{x}}_{p}\in R^{(WH)\times N_{e}}$, where $N_{e}$ is the embedding size. These flattened patches are treated as a sequence of token embeddings ${\rm\bm{z}}_{0}$:
${\rm\bm{z}}_{0}={\rm\bm{x}}=[{\rm\bm{x}}^{1}_{p};...;\bm{x}^{i}_{p};...;\bm{x}^{N}_{p}],$ (1)
where ${\rm\bm{x}}^{i}_{p}$ is the $i$th token embedding in the sequence.
@@
To preserve position information, learnable 2D position embeddings $E_{\mathrm{pos}}$ are added to the patch embeddings. The size of these positional embeddings is fixed and scaled based on the board size and hyperparameters to ensure compatibility with patch embeddings ${\rm\bm{x}}_{0}$. The position embeddings are incorporated into the token embeddings ${\rm\bm{x}}_{0}$ as follows:
${\rm\bm{z}}_{0}\leftarrow{\rm\bm{z}}_{0}+E_{\mathrm{pos}}.$ (2)
The output size of a transformer encoder layer corresponds to the number of input embeddings. For Gomoku, where the action space is $HW$, AlphaViT requires $HW$ embeddings. To achieve this, we set $k=2n+1$, where $n$ is a positive integer, stride $s=1$, and padding $pad=\lfloor k/2\rfloor$.
For Othello, the action space is $HW+1$ to accommodate the pass move. Using the same parameters as Gomoku ($k=2n+1$, $s=1$, and $pad=\lfloor k/2\rfloor$), AlphaViT needs one additional embedding for the pass move. To address this, we introduce a learnable pass token $\bm{x}_{\mathrm{pass}}$ and append it to the input embeddings.
To estimate the board value, we prepend a learnable value embedding $\bm{x}_{\mathrm{value}}$. Additionally, to enable AlphaViT to distinguish between different game types, we incorporate a static game embedding $\bm{x}_{\mathrm{game}}$, represented using one-hot encoding. These embeddings are appended to the initial embeddings ${\rm\bm{z}}_{0}$, resulting in
${\rm\bm{z}}_{0}=[\bm{x}_{\mathrm{value}};\bm{x}_{\mathrm{game}};\bm{x}^{1}_{p}\bm{\mathrm{E}};...;\bm{x}^{WH}_{p}\bm{\mathrm{E}};\bm{x}_{\mathrm{pass}}].$ (3)
The position embeddings are added before the pass, value, and game embeddings are appended. This approach allows for scaling of the position embeddings without affecting these embeddings, which do not contain positional information. The resulting sequence ${\rm\bm{z}}_{0}$ serves as the input for the transformer encoder.
@@
Sequence $\bm{z}_{0}$ is fed into the transformer encoder, which consists of $L$ transformer encoder layers. The output of the last encoder layer ${\rm\bm{z}}_{L}$ consists of $M\times N+2$ vectors. The first vector $z^{0}_{L}$ derived from the value embedding is processed by the value head implemented as a multilayer perceptron (MLP) denoted as $\text{MLP}_{v}$. This head estimates the value $v$:
$v=\tanh(\mathrm{NLP}_{v}(LN(z^{0}_{L}))),$ (4)
@@
where LN denotes layer normalization. Tanh activation ensures the value ranges from $-1$ to $1$, representing the winning probability.
@@
The vectors $z^{2}_{L},...,x^{WH+1}_{L}$ derived from the board patches are processed by the policy head, another MLP ($\mathrm{NLP}_{p}$). The policy head outputs the move probability $p_{m,n}$ for each board position $(m,n),0\leq m<H,0\leq n<W$ using the sigmoid function. The policy head outputs matrix $R^{M\times N}$ representing the move probabilities:
$p(m,n)=Sigmoid(\mathrm{MLP}_{p(m,n)}(LN(z^{m*H+n+1}_{L}))).$ (5)
@@
For Othello, $p(m,n-1)$ represents the probability of pass action. For Connect4, since a player can select only $W$ actions, AlphaViT uses only $p(0,n)$.
@@
To decide on a move, AlphaViT employs Monte Carlo Tree Search with Upper Confidence Bound (UCT), using the value and move probability predicted by the DNN. AlphaViT uses the same MCTS algorithm as AlphaZero, as shown in Appendix 1. The parameters used in AlphaViT are presented in Appendix 3.
@@
AlphaViT has a significant drawback: the size of the policy vector is fixed to the input size of the transformer encoder layer. To address this problem, we propose an improved AlphaViT called AlphaViD. Although similar to AlphaViT in many respects, AlphaViD has a transformer decoder, as shown on the left in Fig. 3. AlphaViD uses the transformer encoder and decoder to estimate the value and move probability, respectively.
@@
The input board is linearly embedded using a convolutional layer and fed into a transformer encoder, similar to AlphaViT. The input embedding sequence is as follows
${\rm\bm{z}}_{0}=[\bm{x}_{\mathrm{value}};\bm{x}_{\mathrm{game}};\bm{x}^{1}_{p}\bm{\mathrm{E}};...;\bm{x}^{WH}_{p}\bm{\mathrm{E}}].$ (6)
@@
The embedding composition of AlphaViD is different from that of AlphaViT. Since AlphaViD only estimates the value from the encoder output, it does not require the embedding size to exceed the action space size. The embedding sequence consists of value and game tokens and patch embeddings. The estimated value is obtained using the value head, which processes the output corresponding to the value embedding from the last layer of the transformer encoder.

### v2

# AlphaViT: A Flexible Game-Playing AI for Multiple Games and Variable Board Sizes

@@ Abstract
This paper presents novel game-playing AI agents based on the AlphaZero framework, enhanced with Vision Transformer (ViT): AlphaViT, AlphaViD, and AlphaVDA. These agents are designed to play multiple board games of various sizes using a single network with shared weights, thereby overcoming AlphaZero’s limitation of fixed-board-size constraints. AlphaViT employs only a transformer encoder, whereas AlphaViD and AlphaVDA incorporate both transformer encoders and decoders. In AlphaViD, the decoder processes outputs from the encoder, whereas AlphaVDA uses a learnable embeddings as the decoder input. The additional decoder layers in AlphaViD and AlphaVDA provide flexibility to adapt to various action spaces and board sizes. Experimental results show that the proposed agents, trained on either individual games or multiple games simultaneously, consistently outperform traditional algorithms such as Minimax and Monte Carlo Tree Search and approach the performance of AlphaZero, despite using a single deep neural network (DNN) with shared weights. In particular, AlphaViT shows strong performance across all tested games. Furthermore, fine-tuning the DNN using pre-trained weights from small-board games accelerates convergence and improves performance, particularly in Gomoku. Interestingly, simultaneous training on multiple games yields performance comparable to, or even surpassing, single-game training. These results indicate the potential of transformer-based architectures to develop more flexible and robust game-playing AI agents that excel in multiple games and dynamic environments.
@@ 1 Introduction
Artificial intelligence (AI) has advanced remarkably in recent years, demonstrating its potential across a wide range of applications. One area where AI has excelled is in mastering board games, often outperforming top human players. Notable achievements include AI systems that defeat humans in games, such as checkers, chess (Campbell et al., 2002), and Othello (Buro, 1997). A significant turning point occurred in 2016 when AlphaGo (Silver et al., 2016), an AI designed specifically for the Go game, defeated the world’s top players. After this, AlphaZero (Silver et al., 2018) was introduced, demonstrating its capability to master various board games, including Chess, Shogi, and Go. This further highlights AI’s superhuman skills of AI in this domain.
@@
Despite these successes, many current game-playing AI agents suffer from a fundamental limitation: they are typically designed for only a single specific game and cannot play other games. Even when the game rules remain consistent, these agents cannot handle variations in board size. In contrast, humans can easily switch between different board sizes. For example, beginners in Go often start practicing on smaller boards (e.g., $9\times 9$) before progressing to larger boards (e.g., $19\times 19$). However, AI agents like AlphaZero, which are designed for a single specific game and fixed board size, require substantial reprogramming to accommodate such changes.
@@
For AlphaZero, this limitation arises from its deep neural network (DNN) architecture, which requires a fixed input size. AlphaZero’s DNN consists of residual blocks and multilayer perceptrons (MLPs). These components are designed for a specific input size. The output size of the residual blocks varies with changes in input size, creating inconsistencies with the expected MLP input size. Consequently, AlphaZero cannot function properly even with small variations in board size.
To address this limitation, this study proposes replacing residual blocks in the AlphaZero framework with Vision Transformer (ViT) (Dosovitskiy et al., 2021). ViT is an image-classification DNN based on the transformer architecture. ViT divides an image into patches and infers a class of the image from the patches. A key advantage of ViT is its flexibility in handling images of various sizes. This flexibility enables the AlphaZero framework to adapt to various games and board sizes.
@@
This paper presents game-playing agents capable of handling multiple games and variable board sizes using a single DNN. These agents, named AlphaViT, AlphaViD (AlphaViT with a transformer decoder layer), and AlphaVDA (AlphaViD with learnable action embeddings), are based on the AlphaZero framework. The agents predict the value of a game state and policy using a DNN, while their decisions are made using Monte Carlo Tree Search (MCTS). Our computational experiments show that the proposed agents can be simultaneously trained on three games (Connect4, Gomoku, and Othello) using a single DNN with shared weights. Moreover, the proposed agents outperform traditional algorithms such as Minimax and MCTS across various games while approaching the performance of AlphaZero, whether trained on a single game or simultaneously trained on multiple games.
Portions of this text were previously published as part of a preprint (https://arxiv.org/abs/2408.13871).
@@ 2 Related work
Game-playing AI agents have reached superhuman performance levels in traditional board games such as Checkers (Schaeffer et al., 1993), Othello (Buro, 1997, 2003), and Chess (Campbell, 1999; Hsu, 1999; Campbell et al., 2002). In 2016, AlphaGo (Silver et al., 2016), a Go-playing AI, has defeated the world’s top Go players, marking the first superhuman-level performance in Go. AlphaGo relies on supervised learning from a large database of expert human moves and self-play data. Subsequently, AlphaGo Zero (Silver et al., 2017) has defeated AlphaGo, relying only on self-play data. In 2018, Silver et al. (2018) have proposed AlphaZero, which has no restrictions on playable games. AlphaZero has outperformed other superhuman-level AIs in Go, Shogi, and Chess. Interestingly, AlphaZero’s capabilities extend beyond traditional two-player perfect information games, with research exploring its potential in more complex scenarios. For example, Hsueh et al. (2018) have shown AlphaZero’s potential in nondeterministic games. Other extensions include handling continuous action spaces (Moerland et al., 2018) and support for multiplayer games (Petosa and Balch, 2019). However, AlphaZero cannot play multiple games simultaneously and handle games with the same rules but different board sizes using a single DNN with shared weights. This limitation is due to the use of the DNN in AlphaZero for policy and value estimations. AlphaZero’s DNN consists of residual blocks and MLPs. While the residual blocks can process input images of various sizes, the MLPs fix their input sizes. Consequently, the DNN is not sufficiently flexible to accommodate variations in board size.
Researchers have improved the AlphaZero framework to overcome its limitations. Wu (2019) has tackled some of these limitations by improving the efficiency of AlphaZero-like training in Go. His model introduces techniques such as playout cap randomization and policy target pruning, which significantly accelerate self-play learning. Wu has used global pooling layers to standardize varying board sizes to a fixed size, enabling the model to estimate values using MLPs effectively. Furthermore, Wu has exclusively utilized convolutional layers for policy estimation, thereby avoiding the need for MLPs for policy estimation. This design enables seamless handling of varying board sizes while efficiently calculating the policy. This innovation marks a significant step towards creating more flexible AI agents capable of adapting to different board configurations. Similarly, Soemers et al. (2023) have explored transfer learning across various board games, employing fully convolutional networks with global pooling to enable effective transfer between games with different board sizes, shapes, and rules. Their approach has demonstrated the potential of convolutional architectures, enhanced with global pooling, to generalize across various game scenarios.
@@
To address the limitation of the AlphaZero framework, we focus on the transformer architecture and solve the problem by integrating transformer architecture into the AlphaZero framework. The transformer architecture, initially developed for natural language processing (Vaswani et al., 2017), has shown remarkable effectiveness in various domains. The transformer architecture has also been successfully applied to image-processing tasks. Transformer-based models have achieved exceptional performance in various image-related tasks, including image classification (Dosovitskiy et al., 2021), semantic segmentation (Xie et al., 2021), video classification (Li et al., 2022), and video captioning (Zhao et al., 2022). ViT, introduced by Dosovitskiy et al. (Dosovitskiy et al., 2021), is a remarkable example of a transformer-based model for image processing. ViT has achieved state-of-the-art performance in image classification at the time of its introduction. A key feature of ViT is its independence from the input image size (Dosovitskiy et al., 2021). Unlike convolutional neural networks, which require fixed-size inputs, ViT can process images of various sizes by dividing them into fixed-size patches and treating each patch as a token in the transformer architecture. This flexibility makes ViT highly adaptable and efficient when handling different image sizes. By incorporating the ViT architecture into the AlphaZero framework, we extend its capability to handle various games with different board sizes, enhancing its flexibility.
@@ 3 AlphaViT, AlphaViD, and AlphaVDA
AlphaZero’s game-playing capability is limited to games with specific board sizes and rules used during training. This limitation stems from AlphaZero’s deep neural network (DNN), which consists of residual blocks followed by multilayer perceptrons (MLPs) for value and policy computations. The MLPs receive input from the final residual block and assume a fixed input size. Consequently, AlphaZero’s performance is restricted to games with predetermined board size.
@@
To address this limitation, AlphaViT, AlphaViD, and AlphaVDA are developed as game-playing AI agents based on the AlphaZero framework but using Vision Transformer (ViT) architecture. These agents use a combination of a DNN and MCTS (Fig. 1). The DNN receives the board state, and outputs the value estimates and the move probabilities (policies). The MCTS searches a game tree using the estimated value and move probability. By incorporating ViT instead of residual blocks, AlphaViT, AlphaViD, and AlphaVDA can overcome the limitations of AlphaZero and play games with varying board sizes and rules.
@@
An overview of AlphaViT’s DNN architecture is shown in Fig. 2. The DNN in AlphaViT is based on ViT, which has no input-size limitation and can classify images even if the input image size differs from the training image size. This flexibility enables AlphaViT to play games with different board sizes using the same network. In AlphaViT, the game boards are fed into ViT. Initially, these inputs are transformed into patch embeddings through a convolutional layer. Using the convolution layer allows for easy adjustment of patch division parameters such as patch size, stride, and padding. The final outputs of the encoder layer are used to compute the value and the move probabilities.
@@
AlphaViD and AlphaVDA incorporate both transformer encoder and decoder layers for calculating value and move probabilities, as shown in 3. In AlphaViD, the decoder layer receives the input derived from the output of the encoder layer. In contrast, AlphaVDA employs learnable action embeddings as inputs to the decoder layer. The final outputs of the encoder layer are used to compute the value, while the final outputs of the decoder layer are used to compute the move probabilities.
@@
Importantly, AlphaViT, AlphaViD, and AlphaVDA can play any game that AlphaZero can because they employ the same game-playing algorithm as AlphaZero.
@@
Figure 1: AlphaViT, AlphaViD, and AlphaVDA follow the same procedural framework as AlphaZero. They take the board state and a current player as inputs and determine the next move using Monte Carlo Tree Search (MCTS). MCTS explores the game tree using the value and policy probabilities provided by the DNN.
@@
Figure 2: Overview of the AlphaViT model architecture. An input board state is divided into fixed-size patches and linearly embedded by a convolutional layer. Position embeddings are added to the patch embeddings. The resulting sequence of the patch embeddings is appended with a value embedding, a game embedding, and a pass embedding. The sequence is fed into a transformer encoder. Finally, a multilayer perceptron (MLP) estimates the value and move probabilities by taking the outputs from the last layer of the transformer encoder as its input.
Figure 3: Architecture of AlphaViD and AlphaVDA models. An input board state is divided into fixed-size patches and linearly embedded using a convolutional layer. Position embeddings are added to the resulting patch embeddings to preserve spatial information. The sequence of patch embeddings is appended with a value embedding and a static game embedding. This input sequence is fed into a transformer encoder. The estimated value is obtained from a multilayer perceptron (MLP) that takes the output corresponding to the value embedding from the last layer of the transformer encoder. The output embeddings are input to the transformer decoder. For AlphaViD, the input embeddings for the transformer decoder are generated from the output embeddings corresponding to the board state by a fully connected layer. For AlphaVDM, the input embeddings of the transformer decoder are generated from learnable action embeddings. Finally, the move probabilities are produced by another MLP, which takes the output of the last layer of the transformer decoder.
@@ 3.1 Architectures of DNNs
Given a game state $s$, AlphaViT’s DNN predicts the value $v(s)$ and the move probability distribution $\bm{p}(s)$, which includes $p(a|s)$ for each action $a$. In a board game, $s$ represents the board state, and $a$ denotes a move.
@@
The input to the DNN is an $H\times W\times(2T+1)$ image stack ${\mathbf{x}}\in\mathbb{R}^{H\times W\times(2T+1)}$, consisting of $12T+$ 1 binary feature planes of size $H\times W$. Here, $H$ and $W$ are the dimensions of the board, and $T$ is the number of history planes. The first $T$ feature planes represent the occupancies of the first player’s discs, where a feature value of 1 means that a disc occupies the corresponding cell and, and 0 means it does not. The following $T$ feature planes represent the occupancies of the second player’s discs. The final feature plane represents the color of the current player’s disc, where 1 denotes the first player and -1 denotes the second player.
@@
The convolutional layer divides the image stack ${\mathbf{x}}$ into $P\times P$ patches with stride $s$ and padding $p$, and generates the patch embeddings ${\mathbf{z}}_{0}$. The patch size $P$ corresponds to the kernel size of the convolutional layer. The sequence of patch embeddings ${\mathbf{z}}_{0}$ is expressed as
${\mathbf{z}}_{0}=[{\mathbf{x}}^{1}_{p}{\mathbf{E}};...;{\mathbf{x}}^{i}_{p}{\mathbf{E}};...;{\mathbf{x}}^{N_{p}}_{p}{\mathbf{E}}],$ (1)
where ${\mathbf{x}}^{i}_{p}\in\mathbb{R}^{P^{2}(2T+1)}$ is the $i$th flattened 2D patch, and ${\mathbf{E}}$ is a trainable embedding tensor with the shape $(P^{2}(2T+1),D)$. The kernel of the convolutional layer acts as the tensor ${\mathbf{E}}$, which maps each patch to a $D$-dimensional embedding space. Here, $N_{p}$ is the number of embeddings, and $D$ is the embedding size.
@@
To retain position information, learnable 2D position embeddings ${\mathbf{E}}_{\mathrm{pos}}$ are added to the patch embeddings. These positional embeddings are scaled according to the board size and hyperparameters to match the patch embeddings ${\mathbf{x}}_{0}$. The position embeddings are incorporated into the token embeddings ${\mathbf{z}}_{0}$ as follows:
${\mathbf{z}}_{0}={\mathbf{z}}_{0}+E_{\mathrm{pos}}.$ (2)
The output size of the transformer encoder layer is determined by the number of input embeddings. For Gomoku, where the action space is $HW$, AlphaViT requires $N_{p}=HW$ embeddings. To achieve this, we set the patch size $P=2k+1$, where $k$ is a positive integer, stride $s=1$, and padding $\mathrm{pad}=\lfloor P/2\rfloor$. In contrast, for Othello, the action space is $HW+1$ to include the pass move. Using the same parameters as Gomoku ($P=2k+1$, $s=1$, and $\mathrm{pad}=\lfloor P/2\rfloor$), AlphaViT requires one additional embedding for the pass move. To address this, we introduce a learnable pass token $\bm{x}_{\mathrm{pass}}$ and append it to the input embeddings ${\mathbf{z}}_{0}$. To estimate the board value, we prepend a learnable value embedding $\bm{x}_{\mathrm{value}}$. Additionally, to enable AlphaViT to recognize different game types, we incorporate a static game embedding $\bm{x}_{\mathrm{game}}$, represented using one-hot encoding. These embeddings are appended to the input embeddings ${\mathbf{z}}_{0}$, resulting in
${\mathbf{z}}_{0}=[\bm{x}_{\mathrm{value}};\bm{x}_{\mathrm{game}};\bm{x}^{1}_{p}\bm{\mathrm{E}};...;\bm{x}^{N_{p}}_{p}\bm{\mathrm{E}};\bm{x}_{\mathrm{pass}}].$ (3)
Position embeddings are added before appending the pass, value, and game embeddings. This approach ensures that the position embeddings can be scaled independently without being affected by embeddings that do not inherently contain positional information.
@@
The sequence $\bm{z}_{0}$ is fed into the transformer encoder, which consists of $L$ transformer encoder layers. The output of the final encoder layer ${\mathbf{z}}_{L}$ consists of $M\times N+3$ vectors. The first vector $z^{0}_{L}$ derived from the value embedding is processed by the value head implemented as a multilayer perceptron (MLP) denoted as $\text{MLP}_{v}$. This head estimates the value $v$ using the following equation:
$v=\tanh(\mathrm{NLP}_{v}(\mathrm{LN}(z^{0}_{L}))),$ (4)
@@
where LN represents layer normalization. The tanh activation constrains the value within the range $(-1,1)$, representing the winning probability.
@@
The vectors $z^{2}_{L},...,x^{WH+2}_{L}$, derived from the board patches and the pass embedding, are processed by the policy head, which is implemented as another MLP ($\mathrm{NLP}_{p}$). The policy head outputs the move probability $p(m,n)$ for each board position $(m,n),0\leq m\leq W,0\leq n<H$ using the sigmoid function. The policy head outputs a vector $\mathbb{R}^{H\times W+1}$ representing the move probabilities
$p(m,n)=\mathrm{Sigmoid}(\mathrm{MLP}_{p}(\mathrm{LN}(z^{m*H+n+2}_{L})))(m,n).$ (5)
@@
For Othello, $p(W,0)$ corresponds to the probability of a pass action. In contrast, for Gomoku, a pass action is not used. For Connect4, since players can only select from $W$ actions, AlphaViT uses only $p(m,0)$.
@@
To decide on a move, AlphaViT employs Monte Carlo Tree Search (MCTS) with Upper Confidence Bound (UCT), using the value and move probabilities computed by the DNN. The MCTS algorithm used in AlphaViT is identical to that of AlphaZero, as detailed in Appendix 1. The hyperparameters for AlphaViT are listed in Appendix 3.
@@
AlphaViT’s DNN has a significant drawback: the size of the policy vector is fixed by the input size of the transformer encoder layer. To address this issue, we propose two improved DNN architectures for AlphaViD and AlphaVDA. Although their DNNs share many similarities with AlphaViT’s DNN, they employ both a transformer encoder and a transformer decoder to calculate the value and move probability.
@@
In AlphaViD and AlphaVDA, the input board is linearly embedded using a convolutional layer and fed into a transformer encoder, similar to AlphaViT. The input embedding sequence is defined as follows:
${\mathbf{z}}_{0}=[\bm{x}_{\mathrm{value}};\bm{x}_{\mathrm{game}};\bm{x}^{1}_{p}\bm{\mathrm{E}};...;\bm{x}^{N_{p}}_{p}\bm{\mathrm{E}}],$ (6)
@@
where the sequence consists of a value token, a game token, and patch embeddings. The embedding sequence in AlphaViD and AlphaVDA differs from that in AlphaViT. The estimated value is obtained from the value head that processes the output embedding corresponding to the value embedding from the last layer of the transformer encoder. In DNNs of AlphaViD and AlphaVDA, the encoder output is only used directly for value estimation, indicating no need for the embedding size to match the action-space size.